# When Does Divide and Conquer Work for Long Context LLM? A Noise Decomposition Framework

### v1

@@ 6 Conclusion
We presented a theoretical and empirical investigation of divide-and-conquer methods for large context tasks. Our framework separates the final error into task noise (cross-chunk synergy), model noise (per-chunk confusion), and aggregator noise (integration flaws). We showed that chunking can outperform single-shot usage in many tasks where superlinear noise growth appears with context length. However, if cross-chunk synergy is too large or the aggregator prompt is ineffective, chunk-based approaches may fail.
Our experiments illustrate several key behaviors central to our noise decomposition framework, including the sweet-spot trade-off between synergy (task noise) and chunk size (model noise), and the conditions under which weaker models leveraging D&C can outperform stronger single-shot models afflicted by superlinear model noise at large context lengths. We also showed that aggregator noise can be reduced by careful prompt engineering, which can be partially automated using a planner.
In future work, it would be interesting to explore more powerful aggregator models or retrieval-based approaches that reconstruct the relevant cross-chunk information. We hope this framework clarifies the interplay between context length, synergy, and multi-agent design, and provides a path toward improved performance on truly long documents.
Appendix A Justification of Assumption 3.1
Justification by elimination.
We consider different growth rates explicitly:
Sublinear growth.
If $\phi_{\mathrm{model}}(x)$ were sublinear, then the average noise per token would decrease as input length increases, which contradicts empirical observations showing performance degradation with larger contexts.
Linear growth ($O(L)$).
Suppose $\phi_{\mathrm{model}}(x)\approx cL$. Then, dividing the input of length $T$ into $n$ chunks (each of length $T/n$) would yield total model noise:
$\Phi_{\mathrm{model}}=n\cdot c\frac{T}{n}=cT,$
which exactly matches the single-shot model noise. Under strictly linear model noise, chunking would provide no direct benefit in terms of model noise alone.
However, consider the empirical evidence from a key-value (KV) retrieval task: this scenario has negligible task noise ($\Phi_{\mathrm{task}}\approx 0$) and negligible aggregation noise ($\Phi_{\mathrm{agg}}\approx 0$), as retrieval operations are completely independent across chunks. Despite this, we empirically observe a clear and substantial improvement from chunked inference.
Such empirical evidence clearly indicates that linear noise growth cannot fully explain observed performance gains. Therefore, the assumption of linear model noise must be incorrect, and the model noise must instead be strictly superlinear, consistent with polynomial growth ($O(L^{\alpha}),\alpha>1$).
Intermediate growth ($O(L\log L)$).
Suppose $\phi_{\mathrm{model}}(x)\approx cL\log L$. Dividing the input of length $L$ into $n$ chunks of length $L/n$ yields total noise:
$n\cdot c\frac{L}{n}\log\frac{L}{n}=cL(\log L-\log n).$
To examine how small the incremental gain is, consider doubling the number of chunks from $n/2$ to $n$:
$cL(\log L-\log(n/2))-cL(\log L-\log n)=cL(\log n-\log(n/2))=cL\log 2.$
Thus, each time we double the chunk number, we gain only a constant additive improvement ($cL\log 2$), which is independent of $n$. Crucially, this incremental improvement does not increase with larger $n$, but remains constant.
Empirically, however, we observe significant and increasing performance improvements as we continue to subdivide the context into smaller chunks (as shown in the KV retrieval scenario with negligible task and aggregation noises). Such empirical behavior clearly contradicts the prediction of $O(L\log L)$ complexity, which would provide diminishing returns (constant incremental improvement) upon increasing the chunk number.
Therefore, intermediate complexity growth of the form $O(L\log L)$ cannot adequately explain observed empirical benefits, indicating that the true complexity must be strictly faster (e.g., polynomial superlinear growth).
Polynomial superlinear growth ($O(L^{\alpha})$, $\alpha>1$).
Under polynomial superlinear growth, dividing input into $n$ chunks yields total noise:
$n\cdot c\left(\frac{L}{n}\right)^{\alpha}=cn^{1-\alpha}L^{\alpha}.$
Since $\alpha>1$, the factor $n^{1-\alpha}$ decreases significantly with increasing $n$, aligning closely with empirical observations of substantial improvements.
Exponential growth ($O(\exp(L))$).
Exponential complexity would quickly make inference infeasible for even moderate lengths, clearly inconsistent with the practical performance of real-world language models.
Other intermediate or exotic complexities.
While other theoretical complexities (e.g., $O(L\log^{k}L)$, $O(L^{\log L})$, $O(\exp(\sqrt{L}))$) exist, they either (1) fail to match empirical performance characteristics (like offering insufficient improvement), or (2) represent unrealistic scenarios never observed empirically in LLM inference tasks.
Thus, by explicit elimination, empirical evidence strongly supports polynomial superlinear complexity $\phi_{\mathrm{model}}(x)\geq cL^{\alpha}$ with $\alpha>1$. ∎
Appendix B Proof of Thm 3.2
Proof.
Step 1. Bounding $\Phi_{\mathrm{model}}$ via superlinear local noise reduction. Recall the definition of the aggregate model noise:
$\Phi_{\mathrm{model}}=\prod_{i=1}^{n}\Bigl{(}f_{i}^{*}(x_{i})+\phi_{\mathrm{model}}(x_{i})\Bigr{)}-\prod_{i=1}^{n}f_{i}^{*}(x_{i}).$
This term captures the noise introduced by the worker models $\phi_{\mathrm{model}}(x_{i})$. By Assumption 3.1, we assume the local model noise $\phi_{\mathrm{model}}(L)$ grows superlinearly with chunk length $L$, specifically $|\phi_{\mathrm{model}}(L)|\geq cL^{\alpha}$ for constants $c>0$, $\alpha>1$, and sufficiently large $L$.
When we partition the input $x$ of length $T$ into $n$ chunks $x_{i}$, each of length approximately $T/n$, the magnitude of the local noise $|\phi_{\mathrm{model}}(x_{i})|$ is significantly reduced compared to the single-shot noise $|\phi_{\mathrm{model}}(T)|\geq cT^{\alpha}$. While the exact form of $\Phi_{\mathrm{model}}$ involves combinations of these local noise terms (potentially multiplied by ideal output terms $f_{j}^{*}$), the dominant factor driving the change in the magnitude $|\Phi_{\mathrm{model}}|$ is this substantial, superlinear reduction in the individual $|\phi_{\mathrm{model}}(x_{i})|$ terms achieved through chunking.
We can characterize the scaling of $|\Phi_{\mathrm{model}}|$ by considering an upper bound proportional to the sum of the bounds on these dominant local noise contributions. Specifically, when local noises $\phi_{\mathrm{model}}(x_{i})$ are small, the first-order terms in the expansion of $\Phi_{\mathrm{model}}$ involve a sum over terms linear in each $\phi_{\mathrm{model}}(x_{i})$ (e.g., $\sum_{i}(\phi_{\mathrm{model}}(x_{i})\prod_{j\neq i}f_{j}^{*}(x_{j}))$ if $\prod$ is multiplication). The overall magnitude $|\Phi_{\mathrm{model}}|$ can thus be bounded by a term related to the sum of the local noise bounds:
$|\Phi_{\mathrm{model}}|\leq B\cdot\sum_{i=1}^{n}|\phi_{\mathrm{model}}(x_{i})|\leq B\cdot n\cdot\sup_{i}|\phi_{\mathrm{model}}(x_{i})|$
where B accounts for the influence of the $f^{*}_{j}$ terms and higher-order noise interactions (assumed reasonably bounded). Applying the local superlinear bound $|\phi_{\mathrm{model}}(L)|\leq c^{\prime}L^{\alpha}$ (using $c^{\prime}$ for an upper bound constant potentially different from $c$):
$|\Phi_{\mathrm{model}}|\leq B\cdot n\cdot c^{\prime}\left(\frac{T}{n}\right)^{\alpha}=(Bc^{\prime})\cdot n^{1-\alpha}T^{\alpha}.$
Let $C_{bound}=Bc^{\prime}$. Since $\alpha>1$, we have $01-\alpha<$ 0. Therefore, increasing the number of chunks $n$ decreases the factor $n^{1-\alpha}$, ensuring that this upper bound on $|\Phi_{\mathrm{model}}|$ grows slower than $T^{\alpha}$. Specifically, for sufficiently large $T$, we can choose $n$ such that:
$|\Phi_{\mathrm{model}}|\leq C_{bound}\,n^{1-\alpha}T^{\alpha}<cT^{\alpha}\leq|\phi_{\mathrm{model}}(T)|.$
Thus, the total model noise introduced by the D&C approach, $|\Phi_{\mathrm{model}}|$, can be made strictly smaller than the noise magnitude of the single-shot approach, $|\phi_{\mathrm{model}}(T)|$, by appropriately choosing the number of chunks $n$.
Step 2. Bounded mismatch $\Delta^{*}(x)$ (Assumption A2). We explicitly assume there exists a constant $\Delta_{0}$ such that:
$|\Delta^{*}(x)|=\left|\prod_{i=1}^{n}f_{i}^{*}(x_{i})-\prod_{i=1}^{n}g^{*}(x_{i})\right|\leq\Delta_{0},$
independent of the input length $T$ and chunking strategy. Thus, $\Delta^{*}(x)$ does not scale with input size.
Step 3. Bounded aggregator noise $\Phi_{\mathrm{agg}}$ (Assumption A3). We explicitly assume there exist constants $M_{\mathrm{agg}}>0$ and $n_{\mathrm{agg}}>0$ such that:
$|\Phi_{\mathrm{agg}}|\leq M_{\mathrm{agg}},\quad\text{for }n\leq n_{\mathrm{agg}}.$
This ensures aggregation noise remains bounded for a controlled number of chunks.
Step 4. Bounded cross-chunk mismatch $\Phi_{\mathrm{task}}$ (Assumption A4). We explicitly assume there exist constants $M_{\mathrm{task}}>0$ and $n_{\mathrm{task}}>0$ such that:
$|\Phi_{\mathrm{task}}|=|\phi_{\mathrm{task}}(x_{1},\dots,x_{n})|\leq M_{\mathrm{task}},\quad\text{for }n\leq n_{\mathrm{task}}.$
This ensures task noise does not grow significantly when using sufficiently few chunks.
Step 5. Combining all terms. By combining Steps 1–4, the total error of the D&C approach is:
$|f_{\mathrm{D\&C}}(x;n)-g(x)|$ $\leq|\Delta^{*}(x)|+|\Phi_{\mathrm{model}}|+|\Phi_{\mathrm{agg}}|+|\Phi_{\mathrm{task}}|$
$\leq\Delta_{0}+|\Phi_{\mathrm{model}}|+M_{\mathrm{agg}}+M_{\mathrm{task}}.$
Using the bound derived in Step 1, for $n$ chosen appropriately (e.g., $n\approx T/L_{0}$, subject to $n\leq\min(n_{\mathrm{agg}},n_{\mathrm{task}})$):
$|f_{\mathrm{D\&C}}(x;n)-g(x)|\leq\Delta_{0}+(C_{bound}\,n^{1-\alpha}T^{\alpha})+M_{\mathrm{agg}}+M_{\mathrm{task}}.$
Since $\alpha>1$, the term $C_{bound}\,n^{1-\alpha}T^{\alpha}$ grows slower than $T^{\alpha}$. The terms $\Delta_{0},M_{\mathrm{agg}},M_{\mathrm{task}}$ are constants. The single-shot noise magnitude $|\phi_{\mathrm{model}}(T)|$ grows at least as fast as $cT^{\alpha}$. Therefore, for sufficiently large $T$, the sum of the bounded terms and the slower-growing model noise term will be less than the rapidly growing single-shot noise:
$\Delta_{0}+(C_{bound}\,n^{1-\alpha}T^{\alpha})+M_{\mathrm{agg}}+M_{\mathrm{task}}<cT^{\alpha}\leq|\phi_{\mathrm{model}}(T)|.$
Thus,
$|f_{\mathrm{D\&C}}(x;n)-g(x)|<|\phi_{\mathrm{model}}(T)|,$
demonstrating that the divide-and-conquer inference error can be made smaller than the single-shot model noise magnitude for sufficiently large $T$. ∎
Appendix C Sample prompts for planner, manager and worker agents
@@
C.1 Example: Summerization
To further illustrate the planner, we use summerization task as another example. We first show the manual (default) prompts for worker and manager agents as below.
@@ C.3 Example: Math
Appendix D Model noise results
This section provides the detailed numerical data that form the basis for Figure 2 presented in Section 5.2 of the main paper. Figure 2 illustrates the performance degradation of single-agent Large Language Models (LLMs) as the input context length increases, a phenomenon we attribute to model noise ($\phi_{model}(x)$). Table 1 tabulates the specific accuracy scores for the evaluated models (gpt4o, gpt4omini, llama3b, and llama70b) on the Key-Value Retrieval (KV) and Math Find Number (Math) tasks, across a range of input lengths from 1K to 128K tokens. These raw scores directly support the visualizations and analysis concerning the growth of model noise with increasing context size discussed in the main text.
@@ Appendix D Model noise results
Appendix E Task noise results
This section contains the detailed numerical results corresponding to Figure 3 in Section 5.3 of the main paper. Figure 3 illustrates the joint effect of task noise ($\Phi_{task}$) and model noise ($\Phi_{model}$), showcasing how performance varies for different LLM agents on 128K-token length tasks when the effective model context length (i.e., chunk size in the D&C setting) is altered. Table 2 lists the precise performance metrics (accuracy, F1 score, or ROUGE score, depending on the task) for the models gpt4omini, llama70b, and qwen72b. These scores are provided for all six benchmark tasks (KV, Math, QA-IB, QA-LB, Sum, Char) across various model context lengths ranging from 1K to 64K tokens (when processing a total input of 128K tokens). This data underpins the analysis of the three noise regimes and the conditions under which D&C approaches are advantageous, as discussed in the main text.
@@ Appendix E Task noise results
Appendix F Framework utility of predictive parameter estimation
A important measure of a theoretical framework’s value lies in its empirical utility, extending beyond explanatory power to offer practical benefits. This section discusses both the explanatory strengths of our noise decomposition framework and its predictive utility in efficiently determining optimal parameters for Divide and Conquer (D&C) strategies.
F.1 Explanatory Power
Our framework provides a principled understanding of why and when D&C methods succeed or fail in the context of long-document processing by Large Language Models (LLMs). By decomposing the overall error into distinct model noise ($\Phi_{model}$), task noise ($\Phi_{task}$), and aggregator noise ($\Phi_{agg}$), the framework facilitates the disentanglement of various contributing factors to performance. This structured understanding has been instrumental in deriving the key insights presented throughout this paper, such as the conditions favoring chunking and the impact of superlinear model noise growth.
F.2 Predictive Utility for Optimal Chunk Size Estimation
A practical challenge in deploying D&C methods is the selection of an optimal chunk size, which often requires costly exhaustive grid searches over various configurations. We conducted new experiments to evaluate whether our framework’s insights could guide a more efficient estimation of this parameter, particularly in regimes dominated by model noise (Case 3 in Section 3.4), where chunk size significantly impacts performance. In other regimes (Case 1 and Case 2), performance tends to be less sensitive to chunk size variations, making this estimation less critical.

### v2

@@ 6 Conclusion
In this work, we presented a theoretical and empirical framework to understand when and why the Divide-and-Conquer (D&C) strategy is effective for long-context LLMs. Our framework decomposes long-context failure modes into three fundamental fidelity loss components ($\mathcal{L}_{task},\mathcal{L}_{model},\mathcal{L}_{agg}$) that we intuitively refer to as task noise (cross-chunk synergy), model noise (length-induced confusion), and aggregator noise (integration flaws).
Through our formal analysis and empirical validation, we demonstrated that when the length-induced degradation (model noise) grows superlinearly with context size, chunk-based processing can effectively suppress this loss. Provided that the cross-chunk dependencies (task noise) are manageable, D&C enables a weaker model to significantly outperform a more advanced model operating in a single shot. Furthermore, we introduced an LLM-based Planner to dynamically align worker outputs, demonstrating a practical approach to minimizing aggregator errors.
Overall, by bridging the formal fidelity-loss mechanisms with observable noise patterns, our results highlight a principled and actionable pathway to handling massive contexts in LLMs through carefully managed chunking and aggregation strategies.
Appendix A Usage of LLMs
Large Language Models (LLMs) were minimally employed during manuscript preparation, specifically to enhance the clarity and fluency of the writing. The role of LLMs was strictly limited to minor linguistic refinement, and the authors retain full responsibility for the content and accuracy of the work.
Appendix B Derivation and Approximations
This appendix details the relationship between the exact fidelity product and the additive noise approximation.
Exact Decomposition.
As defined in Section 3, the system fidelity is an exact telescoping product:
$\rho_{\mathrm{sys}}=\rho_{\mathrm{task}}\cdot\rho_{\mathrm{agg}}\cdot\rho_{\mathrm{model}}.$
By defining the log-loss $\mathcal{L}\coloneqq-\log(\rho)$, the decomposition becomes exactly additive in log-space:
$\mathcal{L}_{\mathrm{sys}}=\mathcal{L}_{\mathrm{task}}+\mathcal{L}_{\mathrm{agg}}+\mathcal{L}_{\mathrm{model}}.$
Linear Error Approximation.
For high-fidelity regimes ($\rho\approx 1$), we can recover an intuitive additive error formulation. Let $\epsilon_{\mathrm{task}}\coloneqq 1-\rho_{\mathrm{task}}$, $\epsilon_{\mathrm{agg}}\coloneqq 1-\rho_{\mathrm{agg}}$, and $\epsilon_{\mathrm{model}}\coloneqq 1-\rho_{\mathrm{model}}$. Then
$\rho_{\mathrm{sys}}=\rho_{\mathrm{task}}\rho_{\mathrm{agg}}\rho_{\mathrm{model}}=(1-\epsilon_{\mathrm{task}})(1-\epsilon_{\mathrm{agg}})(1-\epsilon_{\mathrm{model}}).$
Expanding the product and dropping higher-order interaction terms (e.g., $\epsilon_{\mathrm{task}}\epsilon_{\mathrm{agg}}$) yields
$\mathcal{E}_{\mathrm{total}}\coloneqq 1-\rho_{\mathrm{sys}}\approx\epsilon_{\mathrm{task}}+\epsilon_{\mathrm{agg}}+\epsilon_{\mathrm{model}}=(1-\rho_{\mathrm{task}})+(1-\rho_{\mathrm{agg}})+(1-\rho_{\mathrm{model}}).$ (6)
Appendix C Justification of Proposition 3.1
We analyze the crossover point where the D&C strategy outperforms a single strong agent. We derive the asymptotic behavior directly from the Fidelity Decomposition framework.
1. Single Strong Agent (The Brain Fog).
By the Super-Linear Collapse assumption, the strong model suffers from attention dispersion as context grows. Its fidelity loss grows super-linearly:
$\mathcal{L}_{\mathrm{strong}}(T)=\omega(T)\implies\lim_{T\to\infty}\frac{\mathcal{L}_{\mathrm{strong}}(T)}{T}=\infty.$
2. Weak D&C System (Deriving Linearity).
Recall the additive decomposition of the D&C system loss:
$\mathcal{L}_{\mathrm{D\&C}}(T)=\mathcal{L}_{\mathrm{task}}+\mathcal{L}_{\mathrm{agg}}+\mathcal{L}_{\mathrm{model}}.$
We analyze the growth of each term under the Bounded Unit Loss assumption:
  • Model Loss ($\mathcal{L}_{\mathrm{model}}$): The input is split into $n\propto T$ chunks of fixed length. Assuming the error introduced by each local worker is bounded and independent of $T$, the accumulated log-loss scales linearly with the number of chunks: $\mathcal{L}_{\mathrm{model}}(T)=O(n)=O(T)$.
  • Overhead ($\mathcal{L}_{\mathrm{task}}+\mathcal{L}_{\mathrm{agg}}$): Assuming the task is feasible (finite $\mathcal{L}_{\mathrm{task}}$) and the aggregation process scales linearly with the input size (or uses a hierarchical structure with bounded error), the overhead grows at most linearly: $\mathcal{L}_{\mathrm{task}}+\mathcal{L}_{\mathrm{agg}}=O(T)$.
Combining these terms, the total D&C loss is bounded linearly:
$\mathcal{L}_{\mathrm{D\&C}}(T)=O(T).$
3. Asymptotic Crossover.
We compare the growth rates. Since $\mathcal{L}_{\mathrm{strong}}(T)$ grows super-linearly ($\omega(T)$) while $\mathcal{L}_{\mathrm{D\&C}}(T)$ grows linearly ($O(T)$), the ratio diverges:
$\lim_{T\to\infty}\frac{\mathcal{L}_{\mathrm{strong}}(T)}{\mathcal{L}_{\mathrm{D\&C}}(T)}=\infty.$
This shows that there exists a threshold $T_{0}$ such that for all $T>T_{0}$, $\mathcal{L}_{\mathrm{D\&C}}<\mathcal{L}_{\mathrm{strong}}$.
Appendix D Tasks description
We experiment on six diverse tasks including: Key-Value Retrieval, Math Find Number, Summarization, Dialogue Character Inference, and Open Question QA with and without choices. These tasks are based on InfiniteBench Zhang et al. (2024b) and LongBench-V2 Bai et al. (2024) but we have modified the generation and prepared different lengths of these tasks. These tasks include: Key-Value Retrieval (KV) We randomly generated Key-Value pairs. For a given key, the task is to retrieve the value associated. This task only evaluates the capability of retrieval. The evaluation metric is accuracy. We prepare synthetic KV task to have lengths ranging from 1K to 128K. Math Find Number (Math) We randomly generated a long list of integers following Gaussian distribution and the task is to find the Kth largest or smallest number in this list. Each query is slightly different in K. This task evaluates the memory and mathematical reasoning. The evaluation metric is accuracy. We prepare synthetic Math task to have lengths ranging from 1K to 128K. Summarization (Sum) We provide English texts and the task is to summarize them. The ground truth summarization comes from the InfiniteBench. This task evaluates the language summarization. The evaluation metric is ROUGE score. Open Question QA (QA-IB and QA-LB) We provide English text and some open questions for question answering. These questions are very diverse and the task requires LLM to do multi-hop reasoning to answers. The task QA-IB comes from InfiniteBench and doesn’t come with choices and the task QA-LB comes from LongBench-V2 and is provided with four choices. The evaluation metrics are respectively F1 score and accuracy. Dialogue Character Inference (Char) We provide English dialogue scripts between many characters. One of the characters is masked and the task is to infer the name of the masked person. This task evaluates the capacity of language reasoning and will require LLM to focus on the interaction of people discussions to understand the relation. The evaluation metric is accuracy. This task comes from InfiniteBench.
Appendix E Sample prompts for planner, manager and worker agents
@@
E.1 Example: Summarization
To further illustrate the planner, we use Summarization task as another example. We first show the manual (default) prompts for worker and manager agents as below.
@@ E.3 Example: Math
Appendix F Model noise results
This section provides the detailed numerical data that form the basis for Figure 2 presented in Section 5.2 of the main paper. Figure 2 illustrates the performance degradation of single-agent Large Language Models (LLMs) as the input context length increases, a phenomenon consistent with length-induced model degradation. Table 2 tabulates the specific accuracy scores for the evaluated models (gpt4o, gpt4omini, llama3b, and llama70b) on the Key-Value Retrieval (KV) and Math Find Number (Math) tasks, across a range of input lengths from 1K to 128K tokens. These raw scores directly support the visualizations and analysis concerning the growth of model error with increasing context size discussed in the main text.
@@ Appendix F Model noise results
Appendix G Task noise results
This section contains the detailed numerical results corresponding to Figure 3 in Section 5.3 of the main paper. Figure 3 illustrates the joint effect of cross-chunk dependency (the task term, $\mathcal{L}_{\mathrm{task}}$) and length-induced degradation (the model term, $\mathcal{L}_{\mathrm{model}}$), showcasing how performance varies for different LLM agents on 128K-token length tasks when the effective model context length (i.e., chunk size in the D&C setting) is altered. Table 3 lists the precise performance metrics (accuracy, F1 score, or ROUGE score, depending on the task) for the models gpt4omini, llama70b, and qwen72b. These scores are provided for all six benchmark tasks (KV, Math, QA-IB, QA-LB, Sum, Char) across various model context lengths ranging from 1K to 64K tokens (when processing a total input of 128K tokens). This data underpins the analysis of the three noise regimes and the conditions under which D&C approaches are advantageous, as discussed in the main text.
@@ Appendix G Task noise results
Appendix H Framework utility of predictive parameter estimation
An important measure of a theoretical framework’s value lies in its empirical utility, extending beyond explanatory power to offer practical benefits. This section discusses both the explanatory strengths of our noise decomposition framework and its predictive utility in efficiently determining optimal parameters for Divide and Conquer (D&C) strategies.
H.1 Explanatory Power
Our framework provides a principled understanding of why and when D&C methods succeed or fail in the context of long-document processing by Large Language Models (LLMs). By decomposing the overall error into distinct terms—task ($\mathcal{L}_{\mathrm{task}}$), model ($\mathcal{L}_{\mathrm{model}}$), and aggregator ($\mathcal{L}_{\mathrm{agg}}$)—the framework facilitates the disentanglement of various contributing factors to performance. This structured understanding has been instrumental in deriving the key insights presented throughout this paper, such as the conditions favoring chunking and the impact of accelerated performance degradation with length.
H.2 Predictive Utility for Optimal Chunk Size Estimation
A practical challenge in deploying D&C methods is the selection of an optimal chunk size, which often requires costly exhaustive grid searches over various configurations. We conducted new experiments to evaluate whether our framework’s insights could guide a more efficient estimation of this parameter, particularly in regimes dominated by model noise (Regime 3 in Section 3.6), where chunk size significantly impacts performance. In other regimes (Regime 1 and Regime 2), performance tends to be less sensitive to chunk size variations, making this estimation less critical.